OpenStudy (anonymous):

Input in standard form the equation of the given line. The line with m = -3/8 and passing through (5, -4)

OpenStudy (saifoo.khan):

Lol. Stay Glad. ;)

OpenStudy (anonymous):

Ready to work on this together?

OpenStudy (saifoo.khan):

@uber.. i aint used that, Lol.

OpenStudy (saifoo.khan):

Haha!

OpenStudy (anonymous):

@Myhuyck: Lemme know if you want to work this out step by step.

OpenStudy (anonymous):

saifoo will be sensei before august 7. bets?

OpenStudy (anonymous):

Probably. I could probably get there faster by just giving people answers instead of working with them step by step. :P

OpenStudy (saifoo.khan):

Lol, thnks for that Satealite! :)

OpenStudy (anonymous):

saifoo. focus haha

OpenStudy (anonymous):

Sure, that's it.

OpenStudy (anonymous):

Ready, mlhuyck?

OpenStudy (saifoo.khan):

Uber, JEALOUSY!

OpenStudy (saifoo.khan):

Mlhy, lol, tryin to..

OpenStudy (anonymous):

First off, do you know the formula, y-y1 = m(x-x1) ?

OpenStudy (saifoo.khan):

brb in 3mins! :/

OpenStudy (anonymous):

\[y-y_1=m(x=x_1)\] with \[m=-\frac{3}{8}, x_1=5,y_1=-4\]

OpenStudy (anonymous):

let me be quiet. you got it

OpenStudy (anonymous):

@Saifoo: Honestly, getting a few medals is not that important to me. I'm more happy knowing that I'm helping folks actually learn and do well on their exams. Knowing that I've got folks succeeding, even if it's a little bit, means far more to me than a few points on some unknown site. :P Oh, and satellite, no need to be quiet lol. I get the impression that mlhuyck is ignoring me anyway, so you might as well as just give her the steps before someone else just posts that the answer is 3x+8y = -17